If ST equals 13 find SQ
1
Expert's answer
2014-11-18T10:14:47-0500
If QRST is a square and ST = 13, then SQ is a diagonal of this square and is also a hypotenuse of the right triangle QST, in which QT = ST, so SQ = (13^2 + 13^2)^0.5 = 13*(2)^0.5 = 18.39.